# Multimodal Fusion and Vision-Language Models: A Survey for Robot Vision

### v1

@@ 5.2 Performance comparison on 3D Object Detection
On the nuScenes benchmark (Table 6), various 3D object detection methods exhibit different advantages and limitations under different sensor fusion strategies. Single-modality methods, such as LiDAR-only [254, 13, 255] or camera-only [256, 257, 258], perform well in specific tasks but remain constrained in complex environments. In contrast, multimodal fusion methods effectively combine LiDAR, camera, and radar, significantly improving detection accuracy, robustness, and environmental adaptability. LiDAR-camera fusion leverages the precise geometric information of LiDAR and the rich semantic features of cameras, excelling in object localization and classification. AutoAlignV2 [259] enhances cross-modal feature fusion through deformable feature alignment, making it more effective in small-object detection (e.g., pedestrians and bicycles). SimpleBEV [260], adopting a BEV-based multi-scale fusion strategy, strengthens depth perception and maintains stable overall detection performance. However, LiDAR-camera fusion methods are susceptible to low-light conditions and adverse weather, particularly when LiDAR sampling density decreases, leading to potential information loss. In contrast, radar-camera fusion enhances motion detection by leveraging radar velocity information, with RCBEVDet++ [263] demonstrating greater stability in dynamic scenes, such as high-speed driving and urban intersections. Furthermore, since radar is not affected by lighting conditions or weather, this fusion strategy exhibits higher robustness in low-visibility environments. However, due to radar’s lower point cloud resolution, its performance in static object detection (e.g., parked vehicles and road signs) still lags behind LiDAR-camera fusion methods. Recently, Transformer-based fusion methods, such as BEVFusion [88] and TransFusion [13], have optimized cross-modal information interaction through adaptive feature alignment, achieving superior performance in object classification and detection accuracy. MV2DFusion [262] further surpasses existing methods by overcoming previous limitations. By adopting a modality-specific query-based fusion mechanism, it enhances dynamic object detection, outperforming all prior methods in vehicle, pedestrian, and bicycle detection, setting a new state-of-the-art (SOTA) performance. Overall, MV2DFusion [262] achieves the best performance across multiple object categories, while LiDAR-camera fusion continues to provide high-precision 3D perception, and radar-camera fusion remains more adaptable in challenging environments. Meanwhile, Transformer-driven multimodal fusion methods further optimize cross-modal information integration, offering more intelligent and efficient solutions for autonomous 3D perception.
@@ 5.3 Performance comparison on navigation
Structured planning further refines navigation strategies. HAMT [120] leverages history-aware transformers for long-horizon reasoning, while TD-STP [121] and DUET [115] employ structured transformer-based planning for better global search efficiency. BEVBert [116], integrating hybrid topo-metric pretraining, achieves state-of-the-art SR and SPL, demonstrating superior generalization across seen and unseen environments.
@@ 5.4 Performance comparison on 3D Question Answering and Scene Understanding
Table 8: Performance comparison of state-of-the-art methods evaluated on the ScanQA validation split. Metrics include CIDEr (C), BLEU-4 (B-4), METEOR (M), and ROUGE (R). Higher values indicate superior performance.
@@
In the ScanQA benchmark (Table 8), various Vision-Language Models (VLMs) exhibit different strengths and limitations across key metrics such as CIDEr, BLEU-4, METEOR, and ROUGE. These methods enhance 3D scene understanding through multimodal learning, leveraging spatial reasoning, cross-modal alignment, and language generation techniques to improve question answering (QA) and scene description tasks.Among them, LLaVA-3D [280] and 3D-LLM [282] demonstrate strong generalization capabilities, excelling in question-answering accuracy and spatial reasoning, primarily due to their optimized multi-view feature extraction and 3D spatial embedding mechanisms. 3D-VLP [281] employs a context-aware alignment strategy, performing well in dense scene description tasks, while LL3DA [283] enhances instruction-following accuracy through visual interaction fine-tuning. In comparison, Chat-Scene [286] and Scene-LLM [285] optimize structured embeddings and object-based feature representations, excelling in object-related queries and scene reasoning tasks. Meanwhile, LEO [284], as an embodied intelligence model, exhibits superior adaptability in long-term task planning and interactive tasks. Notably, large-scale pre-trained models such as Qwen2-VL [287] achieve high accuracy in inference-intensive tasks through vision-language alignment, further validating the importance of cross-modal alignment for 3D question answering (3D-QA). Overall, LLaVA-3D and 3D-LLM lead in 3D scene reasoning, Chat-Scene and Scene-LLM perform best in object-based understanding tasks, and LEO excels in long-term task execution, reflecting the trade-offs between generalization capability, spatial alignment, and embodied interaction. Future research should further optimize multimodal fusion strategies to enhance the adaptability of 3D vision-language models in real-world 3D scenes.
@@ 6.1 Low-quality Data
Incomplete multimodal learning presents challenges in fusion tasks, as missing modalities due to device issues, data loss, or user preference affect performance. For instance, in medical diagnosis, key modalities like PET scans may be unavailable. Solutions fall into imputation-based [302] and imputation-free methods [303]. Imputation-based methods recover missing data via matrix decomposition, graph learning, or GANs, while imputation-free approaches leverage available modalities through latent representation learning and deep learning. Challenges remain in assessment, uneven modality information, and incomplete annotations. Modality imbalance also affects performance, including property and quality discrepancies, causing models to over-rely on certain modalities. To mitigate this, optimization techniques, gradient modulation, and adaptive architectures [304, 305] help balance training. Additionally, knowledge distillation, contrastive learning, and dynamic data augmentation enhance lower-quality modalities, improving integration and model robustness.
@@ 6.3 Efficient Training and Inference
Efficient training and inference is an important challenge in the practical deployment of multimodal large models of visual language. Current mainstream VLMs [209, 206, 204, 208] are usually pre-trained based on large-scale Transformer architectures, which usually have billions or even tens of billions of parameters, and pre-training on large-scale datasets requires a large amount of computational resources and time costs, which limits the training efficiency and scale-up ability of the models, and makes the model deployment face serious computational bottlenecks in real-world applications. This computational cost not only limits the practical application scope of the model, but also increases energy consumption and cost, especially in edge devices or real-time application scenarios (e.g., autonomous driving). In addition, due to the obvious heterogeneity of visual and linguistic modalities in terms of data dimensions and feature space, the models may face low training efficiency, high inference latency, and high memory consumption during joint optimisation. Therefore, how to design a more lightweight model structure, reduce the computational overhead of cross-modal fusion, and effectively improve the training and inference efficiency are important challenges in current VLMs research. In order to alleviate the above problems, current research mainly proposes solutions from lightweight network architecture design, model compression and knowledge distillation, and dynamic modal fusion. On the one hand, the computational complexity is effectively reduced by introducing lightweight visual language fusion architectures, such as CrossViT [308], MobileViT [231],MoCoViT [309]and other multi-scale Transformer architectures; on the other hand, the knowledge from the large-scale teacher model is migrated to the lightweight student model through knowledge distillation [310], which significantly improves the training efficiency while maintaining the performance of the while significantly improving the training efficiency. In addition, Modality Dropout or Adaptive Modality Selection methods can reduce redundant modal computations. Therefore, how to effectively combine structural design, optimisation strategies and modal fusion mechanisms to achieve efficient training and deployment of large-scale VLMs is an important research topic that needs to be broken through in the field of VLMs.

### v3

@@ 5.2 Performance comparison on 3D Object Detection
On the nuScenes benchmark (Table 7), various 3D object detection methods exhibit different advantages and limitations under different sensor fusion strategies. Single-modality methods, such as LiDAR-only [264, 13, 265] or camera-only [266, 267, 268], perform well in specific tasks but remain constrained in complex environments. In contrast, multimodal fusion methods effectively combine LiDAR, camera, and radar, significantly improving detection accuracy, robustness, and environmental adaptability. LiDAR-camera fusion leverages the precise geometric information of LiDAR and the rich semantic features of cameras, excelling in object localization and classification. AutoAlignV2 [269] enhances cross-modal feature fusion through deformable feature alignment, making it more effective in small-object detection (e.g., pedestrians and bicycles). SimpleBEV [270], adopting a BEV-based multi-scale fusion strategy, strengthens depth perception and maintains stable overall detection performance. However, LiDAR-camera fusion methods are susceptible to low-light conditions and adverse weather, particularly when LiDAR sampling density decreases, leading to potential information loss. In contrast, radar-camera fusion enhances motion detection by leveraging radar velocity information, with RCBEVDet++ [275] demonstrating greater stability in dynamic scenes, such as high-speed driving and urban intersections. Furthermore, since radar is not affected by lighting conditions or weather, this fusion strategy exhibits higher robustness in low-visibility environments.RaCFormer [276] represents a new trend in radar-camera fusion. Its Transformer-based query mechanism can adaptively select and fuse key information from both modalities, making it particularly suitable for detecting moving objects in dynamic environments. Compared to traditional methods such as RCBEVDet++, RaCFormer demonstrates higher robustness and faster responsiveness in challenging scenarios like low-light conditions, high-speed motion, or severe occlusion. However, due to radar’s lower point cloud resolution, its performance in static object detection (e.g., parked vehicles and road signs) still lags behind LiDAR-camera fusion methods. Recently, Transformer-based fusion methods, such as BEVFusion [92] and TransFusion [13], have optimized cross-modal information interaction through adaptive feature alignment, achieving superior performance in object classification and detection accuracy. SparseLIF [272] focuses on sparse representation and long-range perception. By introducing a sparse Transformer architecture for fusing LiDAR and camera features, it effectively models large-scale cross-modal spatial relationships. Compared to methods like BEVFusion [92] or TransFusion [13], SparseLIF significantly reduces computational overhead while maintaining perception accuracy, making it more suitable for real-time autonomous driving scenarios.EA-LSS [273] enhances scene structure understanding in LiDAR-camera fusion frameworks through the integration of depth supervision and multi-scale information fusion mechanisms. Its design emphasizes transferring supervision signals from semantic segmentation to 3D detection, thereby improving the recognition of small and occluded objects. Unlike traditional BEV-based methods such as AutoAlignV2 [269] or SimpleBEV [270], EA-LSS focuses more on reconstructing and fusing deep semantic features, enabling stronger detail recovery and geometric consistency in complex urban semantic perception tasks. MV2DFusion [274] further surpasses existing methods by overcoming previous limitations. By adopting a modality-specific query-based fusion mechanism, it enhances dynamic object detection, outperforming all prior methods in vehicle, pedestrian, and bicycle detection, setting a new state-of-the-art performance.
Recent methods such as SpaRC and RQR3D have introduced novel fusion mechanisms and geometric modeling strategies for multimodal 3D object detection, further advancing the performance ceiling of radar-camera-based approaches. Specifically, SpaRC proposes a spatial-semantic consistency reward strategy that formulates cross-modal feature alignment as a policy learning problem, effectively mitigating semantic drift in fused representations. Compared to representative methods like RaCFormer, SpaRC achieves higher AP on small-object categories, indicating its superior capability in aligning weak semantic cues and modeling fine-grained structures. In contrast, RQR3D focuses on pure vision-based fusion and introduces a recursive query refinement mechanism to dynamically filter and aggregate image features, significantly enhancing the stability of object localization. In terms of overall performance, its mAP reaches 59.7%, surpassing methods such as CRN and HyDra, and demonstrating more balanced detection capabilities for small and medium-sized targets, making it suitable for target perception requirements in high-density urban environments.
Overall, MV2DFusion [274] achieves the best performance across multiple object categories, while LiDAR-camera fusion continues to provide high-precision 3D perception, and radar-camera fusion remains more adaptable in challenging environments. Meanwhile, Transformer-driven multimodal fusion methods further optimize cross-modal information integration, offering more intelligent and efficient solutions for autonomous 3D perception. This trend reflects a broader shift in 3D object detection toward Transformer-based fusion architectures. Compared with traditional convolutional or early fusion strategies, Transformer-based methods excel in their ability to model long-range dependencies and perform adaptive attention over heterogeneous modalities. These capabilities enable more fine-grained and context-aware fusion of semantic, geometric, and motion information, leading to improved robustness and generalization in complex scenes. As the demands for autonomous driving systems continue to increase, Transformer-driven fusion is likely to become the dominant paradigm in future 3D perception frameworks.
@@ 5.3 Performance comparison on navigation
Structured planning further refines navigation strategies. HAMT [124] leverages history-aware transformers for long-horizon reasoning, while TD-STP [125] and DUET [119] employ structured transformer-based planning for better global search efficiency. BEVBert [120], integrating hybrid topo-metric pretraining, achieves state-of-the-art SR and SPL, demonstrating superior generalization across seen and unseen environments.Compared to earlier navigation approaches that mainly rely on sequence modeling, reinforcement learning, or structured planning, recent methods such as Meta-Explore [293], CoNav [294], and SUSA [295] introduce more structured and multimodal strategies, representing the latest advances in embodied navigation. Meta-Explore proposes a hierarchical framework with local goal search and frequency-domain semantic features (Scene Object Spectrum, SOS), effectively correcting navigation errors and improving generalization. CoNav enables model-level collaboration by aligning a 3D-text model with a vision-language navigation agent through cross-modal belief alignment, offering improved spatial-semantic reasoning in ambiguous environments. Among them, SUSA achieves state-of-the-art performance across multiple VLN benchmarks (e.g., R2R, REVERIE, SOON) by constructing hybrid semantic-spatial representations that combine textual semantic panoramas with depth-based exploration maps. This design significantly enhances cross-modal grounding and navigation robustness, making SUSA one of the most competitive and generalizable approaches to date.
Recent methods such as PanoGen++ [296] and VISTA [297] have introduced new breakthroughs in panoramic semantic mapping and robust training strategies, further enhancing the performance of multimodal navigation systems in complex environments.Specifically, PanoGen++ proposes a perception enhancement mechanism based on panoramic semantic graph generation. By performing semantic panorama stitching and contextual modeling of the environment, it enables agents to achieve stronger spatial understanding and instruction grounding in unseen scenarios. On the Room-to-Room dataset under the test unseen setting, PanoGen++ achieves a 75% SR and 64% SPL, significantly outperforming representative approaches such as BEVBert and Meta-Explore, demonstrating superior global semantic mapping and task generalization capabilities.VISTA, on the other hand, introduces a curriculum-style multi-stage training scheme and a multimodal language-vision alignment mechanism. It emphasizes gradually increasing task difficulty during training and integrating temporal-spatial information across modalities, effectively alleviating navigation degradation in unseen environments. Compared to methods such as Meta-Explore, VISTA maintains a 75% SR and 67% SPL in the test unseen setting, surpassing current mainstream SOTA methods and highlighting its strong robustness in path planning and multimodal alignment under unseen environments.
In summary, PanoGen++ and VISTA represent the latest advances in semantic mapping and robust generalization strategies for multimodal navigation. Compared with earlier structured planning methods (e.g., TD-STP) and vision-language pretrained models (e.g., PRET), they achieve comprehensive improvements in navigation accuracy and generalization ability. These methods exemplify the ongoing shift in embodied navigation from ”path learning” to ”semantic understanding-driven” paradigms.
@@ 5.4 Performance comparison on 3D Question Answering and Scene Understanding
Table 9: Performance comparison of state-of-the-art methods evaluated on the ScanQA validation split. Metrics include CIDEr (C), BLEU-4 (B-4), METEOR (M), ROUGE (R) and ExactMatch (E)
. Higher values indicate superior performance.
@@
In the ScanQA benchmark (Table 9), various Vision-Language Models (VLMs) exhibit different strengths and limitations across key metrics such as CIDEr, BLEU-4, METEOR, ROUGE and ExactMatch. These methods enhance 3D scene understanding through multimodal learning, leveraging spatial reasoning, cross-modal alignment, and language generation techniques to improve question answering (QA) and scene description tasks.Among them, LLaVA-3D [224] and 3D-LLM [303] demonstrate strong generalization capabilities, excelling in question-answering accuracy and spatial reasoning, primarily due to their optimized multi-view feature extraction and 3D spatial embedding mechanisms. 3D-VLP [301] employs a context-aware alignment strategy, performing well in dense scene description tasks, while LL3DA [304] enhances instruction-following accuracy through visual interaction fine-tuning. In comparison, Chat-Scene [302] and Scene-LLM [299] optimize structured embeddings and object-based feature representations, excelling in object-related queries and scene reasoning tasks. Meanwhile, LEO [300], as an embodied intelligence model, exhibits superior adaptability in long-term task planning and interactive tasks. Notably, large-scale pre-trained models such as Qwen2-VL [305] achieve high accuracy in inference-intensive tasks through vision-language alignment, further validating the importance of cross-modal alignment for 3D question answering (3D-QA). BridgeQA [298] further advances the state-of-the-art in the field by introducing a hierarchical cross-modal bridging module on top of existing 3D question-answering models. Unlike methods such as LLaVA-3D and 3D-LLM, which primarily rely on multi-view features or spatial embedding optimization, BridgeQA constructs multi-level cross-modal fusion channels between global 3D scene context and fine-grained object features, achieving tighter visual-linguistic alignment and spatial reasoning capabilities. This architectural design effectively addresses the shortcomings of existing methods in complex scenes, where they struggle to integrate global semantic information with local detail information. On multiple core metrics of ScanQA (including ExactMatch, CIDEr, BLEU-4, etc.), BridgeQA achieves state-of-the-art performance, particularly demonstrating higher robustness when handling complex spatial relationships and ambiguous semantic queries, further validating the advantages of its hierarchical bridging strategy in 3D scene understanding. Overall, LLaVA-3D and 3D-LLM lead in 3D scene reasoning, Chat-Scene and Scene-LLM perform best in object-based understanding tasks, and LEO excels in long-term task execution, reflecting the trade-offs between generalization capability, spatial alignment, and embodied interaction. Future research should further optimize multimodal fusion strategies to enhance the adaptability of 3D vision-language models in real-world 3D scenes.
@@ 6.1 Low-quality Data
Incomplete multimodal learning presents challenges in fusion tasks, as missing modalities due to device issues, data loss, or user preference affect performance. For instance, in medical diagnosis, key modalities like PET scans may be unavailable. Solutions fall into imputation-based [320] and imputation-free methods [321]. Imputation-based methods recover missing data via matrix decomposition, graph learning, or GANs, while imputation-free approaches leverage available modalities through latent representation learning and deep learning. Challenges remain in assessment, uneven modality information, and incomplete annotations. Modality imbalance also affects performance, including property and quality discrepancies, causing models to over-rely on certain modalities. To mitigate this, optimization techniques, gradient modulation, and adaptive architectures [322, 323] help balance training. Additionally, knowledge distillation [324], contrastive learning, and dynamic data augmentation enhance lower-quality modalities, improving integration and model robustness.
@@ 6.3 Efficient Training and Inference
Efficient training and inference is an important challenge in the practical deployment of multimodal large models of visual language. Current mainstream VLMs [218, 215, 213, 217] are usually pre-trained based on large-scale Transformer architectures, which usually have billions or even tens of billions of parameters, and pre-training on large-scale datasets requires a large amount of computational resources and time costs, which limits the training efficiency and scale-up ability of the models, and makes the model deployment face serious computational bottlenecks in real-world applications. This computational cost not only limits the practical application scope of the model, but also increases energy consumption and cost, especially in edge devices or real-time application scenarios (e.g., autonomous driving). In addition, due to the obvious heterogeneity of visual and linguistic modalities in terms of data dimensions and feature space, the models may face low training efficiency, high inference latency, and high memory consumption during joint optimisation. Therefore, how to design a more lightweight model structure, reduce the computational overhead of cross-modal fusion, and effectively improve the training and inference efficiency are important challenges in current VLMs research. In order to alleviate the above problems, current research mainly proposes solutions from lightweight network architecture design, model compression and knowledge distillation, and dynamic modal fusion. On the one hand, the computational complexity is effectively reduced by introducing lightweight visual language fusion architectures, such as CrossViT [327], MobileViT [241],MoCoViT [328]and other multi-scale Transformer architectures; on the other hand, the knowledge from the large-scale teacher model is migrated to the lightweight student model through knowledge distillation [329], which significantly improves the training efficiency while maintaining the performance of the while significantly improving the training efficiency. In addition, Modality Dropout or Adaptive Modality Selection methods can reduce redundant modal computations. Therefore, how to effectively combine structural design, optimisation strategies and modal fusion mechanisms to achieve efficient training and deployment of large-scale VLMs is an important research topic that needs to be broken through in the field of VLMs. Despite the remarkable progress of VLMs in cross-modal understanding and reasoning, their application in robotic vision still faces several core challenges. One primary obstacle lies in their limited capability for real-time deployment. Leading VLMs such as CLIP [190], BLIP [223], and GPT-4V [216] are generally constructed with large-scale Transformer architectures, featuring dense self-attention operations and extended token contexts. These result in high computational overhead and inference latency, making such models impractical for embedded systems or edge devices. Moreover, many of these models operate with frozen pretrained parameters, restricting their flexibility in adapting to online tasks.Another critical challenge is the insufficient domain adaptability of current VLMs. Since most models are trained on static, large-scale image-text corpora such as LAION-400M and Conceptual Captions, their alignment mechanisms often struggle to generalize to robotics-specific sensory inputs-such as sparse or low-light scenes, 3D geometry, and embodied interactions. This limitation becomes more pronounced under distribution shifts, including changes in lighting, task formats, or robot embodiment across platforms [330].Promising directions to address them include dynamic model pruning, adapter-based fine-tuning for efficient on-device learning, and reinforcement learning-based grounding mechanisms that enable real-time feedback and task adaptation.
@@ 6.4 Scarce of High-quality Dataset
In addition to limited scale and modality diversity, current multimodal datasets often exhibit cultural and environmental biases that constrain the generalization ability of VLMs in real-world robotic scenarios. Many widely used datasets are built upon Western-centric indoor layouts (e.g., ScanNet [248], Matterport3D [26]), contain only English-language instructions (e.g., Room-to-Room [261] for vision-language navigation), or are collected in urban driving environments (e.g., nuScenes [23], Waymo [246]). As a result, VLMs trained on such data may show reduced robustness when deployed in multilingual, cross-cultural, or rural environments, where object appearances, spatial structures, and human behaviors can differ significantly from those in the training data.
Environmental bias also poses structural challenges to model performance. Most vision-language corpora are collected under favorable conditions-clear weather, adequate lighting, and static scenes-lacking coverage of complex situations such as nighttime, adverse weather, or highly dynamic environments. This data collection bias can lead to severe performance degradation when models are exposed to fog, rain, occlusions, or high-motion scenarios in practical robotic tasks, thereby limiting their robustness and generalization capabilities.Such limitations in generalizability are particularly critical in robotic vision applications, especially when long-term deployment, task transfer, or operation in novel environments is required (e.g., navigation and manipulation). VLMs trained solely on static and homogeneous data sources are thus insufficient for the diverse needs of embodied agents operating in open-world settings.
To mitigate these issues, future research should place greater emphasis on diversity, coverage, and fairness during dataset construction. Promising directions include the development of multimodal datasets that encompass multiple cultures, languages, and environmental variables; the integration of cross-domain training and evaluation protocols to enhance model robustness under distribution shifts; and the adoption of domain adaptation, semantic augmentation, and sim-to-real transfer techniques to improve generalization in complex real-world deployments.
6.5 Perspectives and future directions
Spatial intelligence: As robotic vision systems gradually evolve from static recognition and perception toward active exploration and embodied interaction, spatial intelligence has emerged as a key direction in the next phase of multimodal large model development [334]. Spatial intelligence requires models not only to possess geometric perception capabilities for 3D structures but also to perform memory, reasoning, and prediction over time-particularly maintaining persistent representations of spatial semantics, interaction history, and future actions in dynamic environments [335].Currently, multimodal large models still exhibit limitations in spatial understanding. For example, they often lack the ability to explicitly model the spatial topology of real-world environments, which hinders their capability to perform high-quality causal reasoning across space-language-action domains. To address this, future research should introduce explicit spatial memory mechanisms and structured spatial representations, such as 3D graph structures and topological map embeddings, to better support long-term task context modeling and navigation history encoding. In addition, the spatial projection, path generation, and environmental adaptability of VLMs in 3D environments should be enhanced by incorporating components such as SLAM, egocentric modeling, and visual loop closure, thus forming a spatial cognition closed-loop system tailored for embodied agents. Importantly, spatial intelligence also demands that models possess cross-modal and cross-scale generalization abilities, enabling them to handle a wide range of tasks from micro-level manipulation (e.g., object grasping) to macro-level planning (e.g., indoor pathfinding). Therefore, developing a unified framework that integrates spatial semantic mapping, action sequence prediction, and language-based interaction will be an essential direction for future research in embodied multimodal models. To enhance real-world deployment, further exploration is needed into sparse spatial representations, memory compression, and symbolic abstraction, which can improve task alignment and rapid adaptation in large-scale, unfamiliar environments.
Ethical deployment: With the growing integration of vision-language models (VLMs) into robotic systems, concerns regarding model safety, robustness, and ethical deployment are becoming increasingly prominent. Recent studies have demonstrated that VLMs are susceptible to image perturbations and ambiguous linguistic inputs, potentially leading to incorrect interpretations and unsafe actions in real-world environments [336]. To address these issues, future research should prioritize the development of standardized adversarial evaluation benchmarks that assess the robustness of VLMs under multimodal interference, complex semantics, and domain shifts. In addition, the incorporation of causal reasoning mechanisms and multimodal consistency regularization can significantly improve the explainability and accountability of robotic decision-making. For instance, causal chain-of-thought modeling, supported by datasets such as CELLO [336], may enable robots to generate interpretable and traceable justifications for their actions across different stages of task execution. These approaches can also be extended by designing intervention-aware reasoning modules that allow systems to anticipate and clarify potential failures before action is taken, thereby enhancing overall system transparency and reliability. Moreover, the ethical deployment of VLM-based robotic systems in safety-critical or cross-cultural contexts requires careful attention to human oversight and normative alignment. Recent advances have highlighted the potential of human-in-the-loop trust modeling [337] and ethical instruction tuning [19] to improve the controllability and societal acceptability of robot behavior. In this regard, future systems should include integrated ethical supervision modules capable of recording both attention distributions and decision trajectories, presented through interpretable visual-log interfaces to support auditing and post-hoc analysis. Furthermore, incorporating region-specific ethical guidelines-such as those defined by the IEEE Global Initiative on Ethics of Autonomous and Intelligent Systems or the European Union’s AI Act-into model training and deployment workflows may enhance robots’ cultural adaptability and regulatory compliance. Together, these strategies offer a promising direction toward building vision-language-enabled robotic systems that are not only capable and intelligent, but also trustworthy, transparent, and ethically aligned.
Embodied Cognition and Brain-Inspired Design: Another important direction for the future development of VLM-based robotic systems lies in fostering interdisciplinary collaboration with fields such as cognitive science and neuroscience. Although current VLMs have achieved impressive performance in grounding visual and linguistic information, they still lack the embodied reasoning, memory consolidation, and value-driven decision-making observed in biological agents. Drawing inspiration from studies on the human brain’s multimodal integration-particularly the role of the prefrontal cortex and hippocampus in semantic association and spatiotemporal memory-may provide valuable insights for building more generalizable and adaptive robotic systems. For example, recent advances in computational neuroscience have introduced biologically inspired attention and memory mechanisms that could inform the design of more interpretable and causally coherent vision-language architectures [338]. Moreover, aligning robot learning objectives with models of human goal inference and reward prediction may help VLMs better handle ambiguous, underspecified, or ethically complex scenarios in real-world interactions. Through closer collaboration between robotics, machine learning, and neuroscience, future VLM-augmented robots may move beyond purely data-driven pattern recognition to achieve more grounded, cognitively plausible, and socially aware intelligence.